OpenStudy (anonymous):

When graphing an absolute value inequality, such as -4 ge x le 4 does that mean 2 lines since you have two inequality signs?

OpenStudy (anonymous):

\[-4 \ge x \ge 4 \] rather

OpenStudy (dumbcow):

it depends, do you mean \[-4\le x \le 4\]

OpenStudy (anonymous):

Eh, yes.

OpenStudy (anonymous):

My bad, looking at it and still couldn't get it down right d:P

OpenStudy (dumbcow):

haha ok so x can be any value between -4 and 4. so in this case only one line is needed

OpenStudy (dumbcow):

are you graphin this on a number line

OpenStudy (anonymous):

But it can't be -3 through 3, so it would seem like 2 lines on an x,y graph.

OpenStudy (anonymous):

cartesian, I didn't put the other equation up because I wanted to get the full answer myself :P

OpenStudy (dumbcow):

ok sorry then yes it would be like you have 2 solid lines at x=-4 and x=4 with everything in the middle shaded

OpenStudy (anonymous):

wouldn't it be everything outside shaded? since it can't be -3 to 3, it would be everything else, right?

OpenStudy (anonymous):

oh wait, I get it, I have it backwards

OpenStudy (anonymous):

tyvm for the help :D

OpenStudy (dumbcow):

there you go ;)

OpenStudy (dumbcow):

if the inequalities were switched then you'd be right